# GPFL: A Gradient Projection-Based Client Selection Framework for Efficient Federated Learning

### v1

@@ 4.3 Time Comparison
4.4 Effectiveness of Exploit-Explore
(a) Fixed $\alpha$
(b) Variable $\alpha$
@@ 4.4 Effectiveness of Exploit-Explore
5 Related Work
In Federated learning Li et al. (2023) Chen et al. (2024), client selection algorithms are of paramount importance in determining which clients participate in the training process Fu et al. (2023). Existing client selection methods can be broadly categorized into pre-selection and post-selection methods.
Federated learning pre-selection methods are designed to identify suitable clients for participation before the training process begins. These methods can be broadly classified into pre-selection based on client characteristics and pre-selection based on client performance or reliability. Pre-selection based on client characteristics involves considering factors such as computational capacity, bandwidth, or energy constraints of the clients Xu and Wang (2020) Yu and Li (2021). By evaluating these characteristics, the algorithm can prioritize clients that are well-suited to perform the required computation and communication tasks efficiently. Additionally, client heterogeneity can be considered in the client selection processLuo et al. (2022). Since heterogeneous statistical data can introduce biases during training, which can ultimately result in a degradation of accuracy in FL Wang et al. (2020). Pre-selection based on client performance or reliability could include historical accuracy, update quality, or communication stability Deng et al. (2021) Nishio and Yonetani (2019). Assessing these metrics can help the algorithm favour clients that have demonstrated consistent performance or reliability in previous iterations, increasing the likelihood of obtaining high-quality updates from them.
Post-selection methods are used to choose clients based on their contributions or updates during the training process. These methods operate after the clients have finished their local training and generated updates. There are two common approaches for post-selection: client contribution-based methods and client diversity-based methods. Client contribution-based methods focus on identifying clients whose updates have the most positive impact on the global model. Factors such as the update’s magnitude, its effect on the global model’s performance Nguyen et al. (2020), or the novelty it brings Lin et al. (2022) are typically considered. By prioritizing clients with significant or innovative updates, these methods enhance the overall accuracy and convergence speed of the model. On the other hand, client diversity-based methods aim to select clients that represent various subsets of the client population. This diversity may be based on features, data distribution, or other relevant characteristics Balakrishnan et al. (2022) Tang et al. (2022). Including diverse clients helps mitigate bias and enables the global model to perform well across different client populations. This approach is particularly valuable in FL scenarios characterized by substantial heterogeneity among clients.
6 Conclusion

### v2

@@ VI-D Time Comparison
VI-E Effectiveness of Exploit-Explore
((a)) Fixed $\alpha$
((b)) Variable $\alpha$
@@ VI-E Effectiveness of Exploit-Explore
VII Related Work
In Federated learning [DBLP:journals/tkde/LiWWHWLLH23] [DBLP:journals/csur/ChenZZZY24], client selection algorithms are of paramount importance in determining which clients participate in the training process [fu2023client]. Existing client selection methods can be broadly categorized into pre-selection and post-selection methods.
Federated learning pre-selection methods are designed to identify suitable clients for participation before the training process begins. These methods can be broadly classified into pre-selection based on client characteristics and pre-selection based on client performance or reliability. Pre-selection based on client characteristics involves considering factors such as computational capacity, bandwidth, or energy constraints of the clients [xu2020client] [yu2021toward]. By evaluating these characteristics, the algorithm can prioritize clients that are well-suited to perform the required computation and communication tasks efficiently. Additionally, client heterogeneity can be considered in the client selection process[luo2022tackling]. Since heterogeneous statistical data can introduce biases during training, which can ultimately result in a degradation of accuracy in FL [wang2020optimizing]. Pre-selection based on client performance or reliability could include historical accuracy, update quality, or communication stability [deng2021auction] [nishio2019client]. Assessing these metrics can help the algorithm favour clients that have demonstrated consistent performance or reliability in previous iterations, increasing the likelihood of obtaining high-quality updates from them.
Post-selection methods are used to choose clients based on their contributions or updates during the training process. These methods operate after the clients have finished their local training and generated updates. There are two common approaches for post-selection: client contribution-based methods and client diversity-based methods. Client contribution-based methods focus on identifying clients whose updates have the most positive impact on the global model. Factors such as the update’s magnitude, its effect on the global model’s performance [nguyen2020fast], or the novelty it brings [lin2022contribution] are typically considered. By prioritizing clients with significant or innovative updates, these methods enhance the overall accuracy and convergence speed of the model. On the other hand, client diversity-based methods aim to select clients that represent various subsets of the client population. This diversity may be based on features, data distribution, or other relevant characteristics [balakrishnan2022diverse] [tang2022fedcor]. Including diverse clients helps mitigate bias and enables the global model to perform well across different client populations. This approach is particularly valuable in FL scenarios characterized by substantial heterogeneity among clients.
VIII Conclusion